30 coins having a value of $3.30 consists of nickels, dimes, and quarters. If there are twice as many quarters as dimes, how many coins of each kind were there?
Answer by Alan3354(69443)

30 coins having a value of $3.30 consists of nickels, dimes, and quarters. If there are twice as many quarters as dimes, how many coins of each kind were there?
-------------
N + D + Q = 30
Q = 2D
N + 3D = 30
-----
5N + 10D + 25Q = 330
Q = 2D
5N + 60D = 330
N + 12D = 66
N + 3D = 30
----------------------- Subtract
9D = 36
D = 4 dimes
-----
Sub and find N and Q